OpenStudy (anonymous):

For the following functions, f and g, determine (a) f(g(x)) (b) g(f(x)) If f(x) = 9 x^{3} - 2 x + 11 and g(x) = \sqrt[3]{x}, then (a) f(g(x)) = ; (b) g(f(x)) = .

OpenStudy (nottim):

Ok. So lets start with (a) here.

OpenStudy (anonymous):

ok

OpenStudy (nottim):

f(g(x)) = We see that its f(x) as its "basis"/ the base of the entire equation. And with the info provided, its f(x) = 9 x^{3} - 2 x + 11

OpenStudy (nottim):

So, now theres a g(x) where x would be in f(x)=9 x^{3} - 2 x + 11

OpenStudy (anonymous):

ok

OpenStudy (nottim):

So, x=\sqrt[3]{x} in f(x) = 9 x^{3} - 2 x + 11. Let's sub those in, ok?

OpenStudy (nottim):

Can you do that for me?

OpenStudy (anonymous):

ill try

OpenStudy (nottim):

Once you figure that out, you should be able to continue simplifying, without need of assistance.

OpenStudy (anonymous):

ok i got it

OpenStudy (anonymous):

9x-2x^(1/3)+11 = a and (9x^3-2x+11)^(1/3) = b/ thanks

OpenStudy (nottim):

Uh. That's confusing to read...but all right. IF you need help again, just say it here. I'll come if I can.